# Dual Alignment Maximin Optimization for Offline Model-based RL

### v2

@@ Abstract
Offline reinforcement learning agents face significant deployment challenges due to the synthetic-to-real distribution mismatch. While most prior research has focused on improving the fidelity of synthetic sampling and incorporating off-policy mechanisms, the directly integrated paradigm often fails to ensure consistent policy behavior in biased models and underlying environmental dynamics, which inherently arise from discrepancies between behavior and learning policies. In this paper, we first shift the focus from model reliability to policy discrepancies while optimizing for expected returns, and then self-consistently incorporate synthetic data, deriving a novel actor-critic paradigm, Dual Alignment Maximin Optimization (DAMO). It is a unified framework to ensure both model-environment policy consistency and synthetic and offline data compatibility. The inner minimization performs dual conservative value estimation, aligning policies and trajectories to avoid out-of-distribution states and actions, while the outer maximization ensures that policy improvements remain consistent with inner value estimates. Empirical evaluations demonstrate that DAMO effectively ensures model and policy alignments, achieving competitive performance across diverse benchmark tasks.
Machine Learning, Offline Reinforcement learning
1 Introduction
@@
Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) aims to learn policies from a pre-collected dataset generated by a behavioral policy within the real environment. This paradigm helps avoid the safety risks and high costs associated with direct interactions, making RL applicable in real-life scenarios, such as healthcare decision-making support and autonomous driving (Emerson et al., 2023; Sinha et al., 2022; Mnih et al., 2015). Offline model-based RL (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023) improves upon this by training a dynamics model and using it to generate synthetic data for policy training or planning. The introduction of dynamics models enhances the sample efficiency and allows the agent to answer counterfactual queries (Levine et al., 2020). However, despite strong performance in learned models, policies often degrade significantly when deployed in real environments due to mismatches between synthetic and real distributions.
@@
Most prior works have focused on enhancing the reliability of synthetic samplings from learned models, and directly incorporating off-policy optimization methods, such as SAC (Haarnoja et al., 2018), to address this mismatch. The representative approach is MOPO (Yu et al., 2020), which quantifies model uncertainty by measuring the predicted variance of learned models and subsequently penalizes model-generated trajectories with high uncertainty. This methodology has inspired subsequent research like MOReL (Kidambi et al., 2020) and MOBILE (Sun et al., 2023), which adopt similar uncertainty-aware frameworks. These approaches help generate synthetic data that remains compatible with the offline dataset and construct conservative value functions to mitigate out-of-distribution (OOD) actions. However, creating a perfect learned model from limited offline datasets is impossible, and as illustrated in Fig. 1, simply combining these methods with off-policy mechanisms still fails to resolve the inherent discrepancies in policy behaviors between the learned model and the underlying environment, leading to inconsistent policy performance and leaving the distribution shift problem unsolved.
In this paper, we trace the inherent root of distribution shift challenges in offline RL to the discrepancies between behavior and learning policies in the underlying environmental dynamics. While optimizing for expected returns, we introduce a regularized policy optimization objective for offline RL, that constrains the visitation distribution discrepancies between the behavioral and the learning policies in real environments. We further self-consistently incorporate synthetic data into this objective, deriving a novel maximin optimization objective for offline model-based RL. Building on these objectives, we propose Dual Alignment Maximin Optimization (DAMO), a consistent actor-critic framework that ensures both coherent alignments between model and environment behavior and model-generated and offline data.
@@
DAMO iteratively applies maximin optimization in two phases. In the inner minimization phase, equivalent to critic training, the focus is on the dual conservative value estimation. This value estimation implicitly harmonizes learning policy behaviors in both the model and the environment to avoid OOD states, while explicitly aligning synthetic and offline data to prevent OOD actions. In the outer maximization phase, corresponding to actor learning, policy improvements are guided by the inner value estimates through objective alignment. This preserves reliably conservative behavior and effectively mitigates OOD issues throughout both policy optimization and deployment. Additionally, DAMO incorporates classifiers to approximate the data alignment terms, facilitating practical implementation. Extensive experiments showcase that DAMO successfully mitigates OOD issues and achieves competitive performance across various benchmarks. Our contributions are summarized as the following:
@@
  • We derive a consistent maximin object for offline model-based RL, ensuring both alignments between model and environment policy behaviors, and synthetic and offline data.
@@
  • We develop DAMO, a novel actor-critic framework that employs maximin optimization to achieve the dual conservative value estimation and the consistent policy improvement.
@@
  • We conduct extensive experiments to validate that DAMO effectively alleviates the distribution shift challenge, delivering superior performance across varying benchmark tasks.
@@
Figure 1: The distribution of transition pair $(s,a,s^{\prime})$ across three datasets (Offline Data, Real Data, and Synthetic Data) are shown for the hopper-medium-expert task, comparing two SOTA methods (MOPO and MOBILE) with our proposed DAMO framework. While MOPO and MOBILE align synthetic and offline data well (indicated by a high overlap), they fail to address OOD states in the real environment (circled by an ellipse drawn with a green dotted line). In contrast, DAMO effectively eliminates the OOD region.
2 Related Work
@@
Offline Model-based RL trains a dynamics model using offline data to approximate environmental dynamics and performs conservative policy optimization (Lu et al., 2021) with model-generate data to mitigate OOD issues. MOPO (Yu et al., 2020) penalizes state-action pairs with high model uncertainty, effectively discouraging the agent from selecting OOD actions. MOReL (Kidambi et al., 2020) constructs a pessimistic Markov decision process to prevent the agent from entering OOD regions, ensuring safer and more reliable policy learning. COMBO (Yu et al., 2021) extends Conservative Q-learning (Kumar et al., 2020) to the model-based setting by regularizing the value function on OOD samples. RAMBO (Rigter et al., 2022) employs an adversarial training framework, optimizing both the policy and model jointly to ensure accurate transition predictions while maintaining robustness. MOBILE (Sun et al., 2023) incorporates penalties into value learning by utilizing uncertainty in Bellman function estimates derived from ensemble models. SAMBO (Luo et al., 2024a) imposes reward penalties and fosters exploration by incorporating model and policy shifts inferred through a probabilistic inference framework.
@@
DICE-based Methods. Distribution Correction Estimation (DICE) is a technique that estimates stationary distribution ratios using duality theory. DICE has shown superior performance in evaluating discrepancies between distributions and has been widely applied in various RL domains, including off-policy evaluation (OPE) (Nachum et al., 2019a), offline imitation learning (IL) (Ma et al., 2022), and offline RL (Nachum et al., 2019b). In offline RL, DICE-based methods correct distribution shifts between offline data and environmental dynamics, formulating a tractable maximin optimization objective. For instance, AlgaeDICE (Nachum et al., 2019b) pioneers the application of DICE-based methods in offline RL, employing regularized dual objectives and Lagrangian techniques to address distribution shift challenges. OptiDICE (Lee et al., 2021) incorporates the Bellman flow constraint and directly estimates stationary distribution corrections for the optimal policy, eliminating the need for policy gradients. ODICE (Mao et al., 2024a) combines orthogonal-gradient updates with DICE to enforce state-action-level constraints. Unlike these model-free approaches, DAMO is, to the best of our knowledge, the first method to promote DICE insights in the model-based setting. While model-free methods typically handle discrepancies between two distributions, DAMO skillfully extends this concept by managing the differences across three distributions, using a divergence upper bound to ensure alignment.
Figure 2: Intuiative examples for demystifying OOD issues in offline model-based RL. The behavior policy $\pi_{\beta}$ primarily favors the top road ($a_{1}$) with a minor representation of the bottom road ($a_{3}$) and never selects the middle road ($a_{2}$). Consequently, the dynamics model lacks representation of medium road ($a_{2}$) and treats the $a_{3}$ as an OOD action. The policy $\pi$, trained on both offline and synthetic data, avoids $a_{3}$ when deployed in the real environment but struggles to evaluate $a_{2}$ due to its absence in $M$, resulting in occasionally getting into OOD state. In DAMO, $\pi$ is constrained by $\pi_{\beta}$, guiding the agent to consistently select $a_{1}$. At the bottom, we show that while data alignment ensures a similar distribution between offline and synthetic data, it fails to guarantee behavior consistency between $M$ and $T$.
3 Preliminaries
@@
MDP. We consider the Markov decision process (MDP) defined by the six-element tuple $M=(\mathcal{S},\mathcal{A},T,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $T(s^{\prime}|s,a)$ is the environment transition dynamics, $r(s,a,s^{\prime})>0$ is the reward function, $\mu_{0}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discounted factor. Given an MDP, the objective of RL is to find a policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that maximizes the cumulative reward from the environment. This objective can be formally expressed as: $\pi^{\star}=\arg\max_{\pi}\mathbb{E}_{\mu_{0},\pi,T}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$. In the derivation of DAMO, we utilize the dual form of the RL objective (Puterman, 2014), which is represented as follows:
@@
Here, $\rho^{\pi}_{T}(s,a,s^{\prime})$ is the transition occupancy measure, characterizing the distribution of transition pairs $(s,a,s^{\prime})$ induced by policy $\pi$ under the dynamics $T$, defined by:
$\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}[s_{t}=s,a_{t}=a,\\ s_{t+1}=s^{\prime}|s_{0}\sim\mu_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim T(\cdot|s_{t},a_{t})].$
Here, $\mu_{0}$ denotes the initial state distribution. For simplicity, we use shorthand notation $\rho^{\pi}_{T}$ in the remainder of the article.
Offline Model-based RL. Offline model-based RL aims to find the optimal policy by leveraging a learned dynamics model $M$. Given the offline dataset $\mathcal{D}_{R}$, $M$ is typically trained using maximum likelihood estimation, formulated as $\min_{M}\mathbb{E}_{s,a,s^{\prime}\sim\mathcal{D}_{R}}[-\log M(s^{\prime}|s,a)]$. The training policy $\pi$ will undergo $k$-step rollout in learned dynamics model $M$ and the generated data will be added to a replay buffer $\mathcal{D}_{M}$. Then $\pi$ can be optimized using mini-batches of data sampled from $\mathcal{D}_{R}\cup\mathcal{D}_{M}$, where each datapoint is sampled from the offline dataset $\mathcal{D}_{R}$ with probability $f$ and the model dataset $\mathcal{D}_{M}$ with probability $1-f$. Throughout this work, we assume that the reward function $r(s,a,s^{\prime})$ is known, though it can also be treated as part of the model if unknown.
4 Demystify OOD in Offline Model-based RL
@@
In this section, as illustrated in Fig. 2, we explore the out-of-distribution (OOD) issues in offline model-based RL. We focus on the importance of aligning both the model and offline data, as well as the model and environment policy behavior, which motivates the direction of this research.
4.1 OOD Issues in Offline RL
@@
In offline RL, OOD problems (Mao et al., 2024b) arise when agents encounter states or select actions that fall outside the distribution of the offline data during training or testing.
OOD actions refer to actions that the behavior policy does not choose in specific states, rather than actions that are simply absent from the offline dataset. This implies that OOD actions are inherently state-dependent. Thus some studies use the term OOD state-action pairs to describe the state-action pair $(s,a)$ that does not appear in the offline dataset. Taking OOD actions during training can lead to severely inaccurate value estimation and significantly degraded policy performance in testing. Consequently, avoiding OOD actions has been a focus in previous offline RL research.
@@
OOD states primarily arise in three main scenarios: 1) The learned policy executes unreliable OOD actions, leading to transitions into OOD states. 2) The initial state of the real environment lies outside the offline dataset. Additionally, stochastic dynamics could unexpectedly transition the agent into OOD states, even if it takes in-distribution (ID) actions in ID states. 3) Nonstationary dynamics, such as disturbances in real-world environments, can introduce unexpected changes in state distributions. While this factor is outside the scope of our offline model-based setting, it is an important consideration in real-world applications.
4.2 OOD Issues under Dynamics Model Integration
OOD issues in offline model-based RL are intrinsically linked to the mismatch between synthetic and real data distributions, compounded in model-based settings, where model inaccuracies introduce additional sources of error. More extensive discussions are provided in Appendix B.
OOD Actions. The model, $M$, is used to generate synthetic data for training, but it is essential to avoid taking OOD actions when rolling out policy in the model. This can trigger negative model exploitation (Levine et al., 2020), resulting in poor policy performance. Aligning the distributions of synthetic and offline data can effectively mitigate these issues as it ensures accurate value estimation.
OOD States. The performance of the learning policy during deployment is critically dependent on the alignment of transition pairs between the learned model and the real environment. Significant discrepancies between these transitions can cause performance degradation and lead to OOD states, highlighting the necessity of correcting transition biases during training. Even if the learned policy performs well in the model, it is important, during real-environment deployment, to avoid OOD states, for which we cannot accurately estimate the value function in the offline setting.
5 Dual Alignment Maximin Optimization

### v3

# Dual Alignment Maximin Optimization for Offline MBRL

@@ Abstract.
Offline reinforcement learning agents face significant deployment challenges due to the synthetic-to-real distribution mismatch. While most prior research has focused on improving the fidelity of synthetic sampling and incorporating off-policy mechanisms, the directly integrated paradigm often fails to ensure consistent policy behavior in biased models and underlying environmental dynamics, which arise from discrepancies between behavior and learning policies. In this paper, we first shift the focus from model reliability to constraining policy discrepancies while optimizing for expected returns, and then self-consistently incorporate synthetic data, deriving a novel maximin paradigm, Dual Alignment Maximin Optimization (DAMO). DAMO is a unified framework to ensure both model-environment policy consistency and synthetic and offline data compatibility. The inner minimization performs dual conservative value estimation, aligning policies and trajectories to avoid out-of-distribution states and actions, while the outer maximization ensures that policy improvements remain consistent with inner value estimates. Empirical evaluations demonstrate that DAMO effectively ensures model and policy alignments, achieving competitive performance across diverse offline benchmark tasks.
Key words and phrases:
Offline Model-Based RL, OOD Issues, Maximin Optimization
1. Introduction
@@
Offline reinforcement learning (RL) Lange et al. (2012); Levine et al. (2020) aims to learn policies from a pre-collected dataset generated by a behavior policy within the real environment. This paradigm helps avoid the potential safety risks and high costs associated with direct online interactions, making RL applicable in real-life scenarios, such as healthcare decision-making support and autonomous driving Emerson et al. (2023); Sinha et al. (2022); Mnih et al. (2015). Offline model-based RL (MBRL) Yu et al. (2020); Kidambi et al. (2020); Sun et al. (2023) improves upon this by training a dynamics model and using it to generate synthetic data for policy training. The introduction of the dynamics model enhances the sample efficiency and allows the agent to answer counterfactual queries Levine et al. (2020). However, despite strong performance in dynamics models, trained policies often degrade when deployed in real scenarios due to mismatches between synthetic and real distributions.
@@
Most prior works have focused on enhancing the reliability of synthetic samplings from dynamics models, and directly incorporating off-policy optimization methods such as SAC Haarnoja et al. (2018), to address this mismatch. The representative approach is MOPO Yu et al. (2020), which quantifies dynamics model uncertainty by measuring the predicted variance of learned models and penalizes synthetic trajectories with high uncertainty. This methodology has inspired subsequent researchs like MOReL Kidambi et al. (2020) and MOBILE Sun et al. (2023), which adopt similar uncertainty-aware frameworks. These approaches construct conservative value functions to mitigate overestimation of out-of-distribution (OOD) actions, thereby reducing execution of these unreliable actions and facilitating the generation of synthetic data compatible with the offline dataset. However, as illustrated in Fig. 1, although this uncertainty-aware framework aligns the synthetic data with offline data and mitigates OOD actions during policy training, it fails to alleviate OOD trajectories when deployed into real scenarios, leading to inconsistent policy behavior during deployment and leaving the OOD issues in offline MBRL unresolved.
To bridge this gap, we start from a regularized policy optimization objective, which limits the discrepancy between the visitation distribution of the training policy and the behavior policy under the real environment. By self-consistently incorporating the synthetic data into the regularized objective, we derive a unified maximin optimization objective for offline MBRL. Building on this objective, we propose Dual Alignment Maximin Optimization (DAMO), a unified actor-critic framework to ensure dual conservative estimation.
@@
DAMO iteratively applies maximin optimization in two phases: the inner minimization and the outer maximization. In the inner minimization, equivalent to critic training, we provide a theoretical analysis showing that it enforces dual conservative estimation. This value estimation explicitly aligns the synthetic and offline data to prevent OOD actions during training, while implicitly aligning learning policy behaviors in both the learned model and the real environment to avoid OOD states during deployment. In the outer maximization, corresponding to actor learning, policy improvements are guided by the inner value estimates through objective alignment. This preserves conservative behavior and mitigates OOD issues throughout both policy optimization and deployment. Additionally, DAMO incorporates a classifier to enable the dual conservative estimation during practical implementation. Extensive experiments showcase that DAMO successfully mitigates OOD issues and achieves competitive performance across various offline benchmarks. Our contributions are summarized as follows:
@@
  • We derive a consistent maximin objective for offline MBRL, ensuring both alignments between dynamics model and real environment policy behaviors, and synthetic and offline data.
@@
  • We develop DAMO, an actor-critic framework that employs maximin policy optimization to achieve both the dual conservative value estimation and the consistent policy improvement.
@@
  • We conduct experiments to validate that DAMO alleviates the OOD issues throughout training and deployment, delivering superior performance across varying benchmarks.
@@
Figure 1. The distributions of transition pair $(s,a,s^{\prime})$ across three datasets (Offline Data, Real Data, and Synthetic Data) are shown for the hopper-medium-expert task, comparing MOPO and MOBILE with our proposed DAMO framework. The Real Data is collected through executing the policy trained using specific offline methods in the real environment. While MOPO and MOBILE align synthetic and offline data well (indicated by a high overlap), they fail to mitigate OOD states in the real environment (circled by an ellipse). In contrast, DAMO effectively eliminates the OOD region.
2. Related Work
@@
Offline Model-based RL trains a dynamics model using offline data to approximate environmental dynamics and performs conservative policy optimization Lu et al. (2021) with model-generate data to mitigate OOD issues. MOPO Yu et al. (2020) penalizes state-action pairs with high model uncertainty, effectively discouraging the agent from selecting OOD actions. MOReL Kidambi et al. (2020) constructs a pessimistic Markov decision process to prevent the agent from entering OOD regions, ensuring safer and more reliable policy learning. COMBO Yu et al. (2021) extends Conservative Q-learning Kumar et al. (2020) to the model-based setting by regularizing the value function on OOD samples. RAMBO Rigter et al. (2022) employs an adversarial training framework, optimizing both the policy and model jointly to ensure accurate transition predictions while maintaining robustness. MOBILE Sun et al. (2023) incorporates penalties into value learning by utilizing uncertainty in Bellman function estimates derived from ensemble models. SAMBO Luo et al. (2024b) imposes reward penalties and fosters exploration by incorporating model and policy shifts inferred through a probabilistic inference framework.
@@
DICE-based Methods. Distribution Correction Estimation (DICE) is a technique that estimates stationary distribution ratios using duality theory. DICE has shown superior performance in evaluating discrepancies between distributions and has been widely applied in various RL domains, including off-policy evaluation (OPE) Nachum et al. (2019a), offline imitation learning (IL) Ma et al. (2022), and offline RL Nachum et al. (2019b). In offline RL, DICE-based methods correct distribution shifts between offline data and environmental dynamics, formulating a tractable maximin optimization objective. For instance, AlgaeDICE Nachum et al. (2019b) pioneers the application of DICE-based methods in offline RL, employing a regularized dual objective and Lagrangian techniques to address distribution shift challenges. OptiDICE Lee et al. (2021) incorporates the Bellman flow constraint and directly estimates stationary distribution corrections for the optimal policy, eliminating the need for policy gradients. ODICE Mao et al. (2024b) deploys orthogonalization techniques to eliminate the conflicts between varying gradients, and updates the policy with DICE to enforce state-action-level constraints. Unlike these proposed model-free approaches, DAMO is, to the best of our knowledge, the first method to promote DICE insights in the model-based setting. While model-free methods typically handle discrepancies between two distributions, DAMO skillfully extends this concept by managing the discrepancies across three distributions, using a divergence upper bound to ensure dual alignment.
3. Preliminaries
@@
MDP. We consider the Markov decision process (MDP) defined by the six-element tuple $M=(\mathcal{S},\mathcal{A},T,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $T(s^{\prime}|s,a)$ is the environment transition, $r(s,a,s^{\prime})>0$ is the reward function, $\mu_{0}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discounted factor. Given an MDP, the objective of RL is to find a policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that maximizes the cumulative reward from the environment. This objective can be formally expressed as: $\pi^{\star}=\arg\max_{\pi}\mathbb{E}_{\mu_{0},\pi,T}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$. In this paper, we utilize the dual form of the RL objective Puterman (2014), which is represented as follows:
@@
Here, $\rho^{\pi}_{T}(s,a,s^{\prime})$ is the transition occupancy measure, characterizing the probability distribution of transition pairs $(s,a,s^{\prime})$ induced by policy $\pi$ under the dynamics $T$ among the episode, defined by:
$\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}[s_{t}=s,a_{t}=a,s_{t+1}=s^{\prime}\\ |s_{0}\sim\mu_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim T(\cdot|s_{t},a_{t})].$
$\mathbb{P}$ represents the probability. For simplicity, we will substitute $\rho^{\pi}_{T}(s,a,s^{\prime})$ with notation $\rho^{\pi}_{T}$ in the remainder of the article.
4. Demystify OOD in Offline MBRL
Figure 2. Intuitive examples for the OOD issue. The behavior policy $\pi_{\beta}$ primarily favors the top road ($a_{1}$) with a minor representation of the bottom road ($a_{3}$) and never selects the middle road ($a_{2}$). Thus, the dynamics model lacks representation of the $a_{2}$ and treats the $a_{3}$ as an OOD action. The policy $\pi$, trained on both offline and synthetic data, avoids $a_{3}$ when deployed in the real environment but struggles to evaluate $a_{2}$ due to its absence in $M$, resulting in occasionally getting into OOD state. In DAMO, $\pi$ is constrained by $\pi_{\beta}$, guiding the agent to consistently select $a_{1}$. At the bottom, we show that while data alignment ensures a similar distribution between offline and synthetic data, it fails to guarantee behavior consistency between $M$ and $T$.
@@ 4. Demystify OOD in Offline MBRL
In this section, as illustrated in Fig. 2, we explore the OOD issues in offline MBRL. We focus on the essence of aligning both the synthetic and offline data, as well as the dynamics model and real environment policy behavior, which motivates the direction of this research.
4.1. OOD Issues in Offline RL
@@
In offline RL, OOD problems Mao et al. (2024a) arise when agents encounter states or select actions that fall outside the offline data. Avoiding the OOD issue is a prerequisite for effective policy training.
OOD actions refer to actions that the behavior policy does not choose in specific states, rather than actions that are absent from the offline dataset. The value of OOD actions is prone to overestimation due to limited sample coverage of the offline dataset and the requirement of maximizing the action values, leading to the incorrect or suboptimal direction of policy improvement.
@@
OOD states primarily arise in three main scenarios: 1) The policy executes unreliable OOD actions, leading to transitions into OOD states. 2) Stochastic dynamics could transition the agent into OOD states, even if it takes in-distribution (ID) actions in ID states. Additionally, the initial state of the real environment may lie outside the dataset. 3) Nonstationary dynamics, such as disturbances from real world Luo et al. (2024a) and cross-domain settings Lyu et al. (2025, 2024) can introduce fluctation in state distributions. In this paper, we consider the first two factors. While the third factor is outside the scope of the offline RL setting, it is a critical consideration in some real-world applications.
4.2. OOD Issues under Biased Dynamics Model
The prediction bias and stochasticity of the dynamics model compound the OOD issue in the offline setting. The prediction bias causes a misalignment between synthetic and offline data, and this bias is amplified with iterative use of the model, leading to the undesirable exploitation of OOD state-action pairs Levine et al. (2020). Meanwhile, the stochasticity elevates the likelihood of encountering OOD states, undermining the policy performance across both learned models and real environments. More discussions are given in Appendix B.
OOD actions. The integration of the dynamics model allows the agent to query the possible outcomes of selecting arbitrary actions in any given state. However, the dynamics model cannot provide accurate predictions for OOD actions, and the prediction bias will mislead the estimation of these OOD action values. To address the problem, it is essential to constrain the scope of counterfactual queries within the support of the offline dataset, which amounts to aligning the distribution of synthetic data with the offline data.
OOD states. The agent can mitigate the OOD states stemming from the first scenario by constraining the training policy to avoid OOD actions during training. However, the inherent prediction bias and intrinsic stochasticity of the biased dynamics model will increase the potential of encountering OOD states stemming from the second scenario during deployment. To address this problem, it is essential to ensure that the policy behavior remains consistent across both the dynamics model and the real environment.
In summary, while optimizing the training policy $\pi$ to maximize the expected returns, we should constrain the policy in two directions: 1) aligning the synthetic data $(s,a,s^{\prime})_{M}^{\pi}$ with the offline data $(s,a,s^{\prime})_{T}^{\pi_{\beta}}$, and 2) ensuring that training policy $\pi$ exhibits consistent behaviors in both dynamics models $M$ and real environments $T$.
5. Dual Alignment Maximin Optimization